OpenStudy (anonymous):

The formula F = + 37 estimates the temperature F in degrees Fahrenheit when crickets chirp n times per minute. Use the formula to estimate the temperature when n = 75 chirps/min. Round to the nearest whole number, if necessary.

OpenStudy (lilg132):

i dont understand the question lol

OpenStudy (lilg132):

oh wait

OpenStudy (lilg132):

t = c(0.25) + 37 c = chirps every minute 75 * 0.25 = 18.75 chirps every 15 seconds 18.75 + 37 = 55.75 Fahrenheit

OpenStudy (lilg132):

dont forget to click good answer :)